Review PartA On your first trip to Planet X you happen to take along a 300 g mass, a 40-cm-long spring, a meter stick, and a stopwatch. Youre curious about the free-fall acceleration on Planet X, where ordinary tasks seem easier than on earth, but you cant find this information in your Visitors Guide. One night you suspend the spring from the ceiling in your room and hang the mass from it. You find that the mass stretches the spring by 26.7 cm . You then pull the mass down 11.7 cm and release it. With the stopwatch you find that 9 oscillations take 15.5 Based on this information, what is g? 3x= 2.8 SubmitPre X Incorrect; Try Again; 9 attempts remaining mys evious Ans Reguest Answer < Return to Assignment Provide Feedback
Answer #1

Period

T=15.5/9 =1.722 s

Since

T=2n

4π2 472m x 0.3 T-- 1.7222

SInce force due to gravity is balanced by spring restoring force

mg=Kx

g=Kx/m =3.993*0.267/0.3

g=3.5538 m/s2